# Learning to Match Unpaired Data with Minimum Entropy Coupling

### v1

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
Modeling multimodal data allows for a more comprehensive understanding, reflecting the inherently multi-faceted nature of the real world. Recent works in representation learning (Radford et al., 2021; Lu, 2023; Manzoor et al., 2023; Chen et al., 2023), the study of multivariate systems (Kaplanis et al., 2023; Liang et al., 2023; Bounoua et al., 2024b), generative modeling (Rombach et al., 2022; Tang et al., 2023b, a; Bounoua et al., 2024a; Esser et al., 2024), and multimodal conversational agents (Li et al., 2023; Liu et al., 2023; Shukor et al., 2023; Xue et al., 2024; Wu et al., 2024), are few examples to illustrate the fervent effort in the machine learning community to address and exploit multimodality. However, the intrinsic complexity of multimodal data introduces several challenges that hinder their application in machine learning research. Modality heterogeneity complicates and sometimes impedes geometric comparisons, requiring for example learning a mapping from ambient to latent spaces (Rombach et al., 2022; Tang et al., 2023b; Liu et al., 2023; Bounoua et al., 2024a) or stringent assumptions (Liang et al., 2022; Xia et al., 2023; Dong et al., 2024; Ibrahimi et al., 2024; Zhang et al., 2024b). Alignment across modalities at spatial, temporal or semantic levels is another challenge, which calls for costly pre-processing steps such as synchronization (Hanchate et al., 2024; Chen et al., 2024; Scirè, 2024; Martin-Turrero et al., 2024).
@@
In this work, we study the problem of unpaired multimodal data through the lens of coupling, a fundamental problem in probability theory, that aims at determining the optimal joint distribution of random variables given their marginal distributions, with early attempts at solving it dating back to the work by Fréchet (1951). The pairing problem belongs to a broad class of methods (Den Hollander, 2012; Lin et al., 2014; Benes & Stepán, 2012; Yu & Tan, 2018): some cast it through the lens of information-theoretic quantities, where optimality is defined in terms of Entropy minimization or Mutual Information maximization, others focus on Optimal Transport (OT) (Villani, 2009; Peyré & Cuturi, 2019), where optimality is defined as minimizing the expected value of a transport cost over the joint distribution. Our focus is the Minimum Entropy Coupling (MEC) problem, which aims at finding the joint distribution with the smallest Entropy, given the marginal distribution of some random variables. Recent applications include entropic causal inference (Kocaoglu et al., 2017; Javidian et al., 2021; Compton, 2022), communication systems (Sokota et al., 2022), steganography (de Witt et al., 2022), random number generation (Li, 2021b), dimensionality reduction (Cicalese et al., 2016; Vidyasagar, 2012), lossy compression (Ebrahimi et al., 2024), and multimodal learning (Liang et al., 2024).
While the MEC problem is known to be NP-Hard (Vidyasagar, 2012; Kovačević et al., 2012), the literature contains many approximation and greedy algorithms  (Painsky et al., 2013; Kovacevic et al., 2013; Cicalese et al., 2016; Li, 2021a), and theoretical studies about the approximation qualities of such approaches (Cicalese et al., 2017, 2019). Nevertheless, the vast majority of prior work on the MEC problem focus on discrete distributions: instead, we consider the continuous variant of MEC, and propose a flexible and general solution to the coupling problem for arbitrary, continuous distributions. The MEC problem for continuous random variables is much more complex than its discrete counterpart, and can be ill-defined in certain cases due to the properties of differential Entropy and the challenges inherent to continuous distributions living in an infinite dimensional space.
@@ 2 Problem Formulation
[…]
@@
Early instances of the coupling problem express it through the lenses of OT (Monge, 1781; Kantorovich, 1942). In the simplest (albeit rich and interesting) scenario, the goal is to minimize the transportation cost between distributions $\mathbb{E}_{x,y\sim p^{\theta}_{X,Y}}[||x-y||^{2}]$, with the implicit assumption of $\mathcal{X}=\mathcal{Y}=\mathbb{R}^{N}$ and under the requirement of exact matching (corresponding to $\lambda_{X}=\lambda_{Y}=\infty$). Several interesting extensions, including additional constraints on the joint distribution such as geometry or structural constraints, lead to tailor-made approaches (Villani, 2009; Peyré & Cuturi, 2019). Other than the trivial relaxation of constraints from exact to approximate, a particularly useful extension concerns the entropy-regularized version of this problem, where the cost function is complemented by the entropic term $\mathbb{H}(p^{\theta}_{XY})$. Although MEC is fundamentally different than OT, a link between the two clearly exists. However, a straightforward comparison is not possible, as the entropic term enters the respective minimization problems with different signs. Minimizing $\mathbb{H}(p^{\theta}_{XY})$ directly over other geometric costs (like the euclidean norm considered in OT) has several advantages in terms of generality, as it does not require geometrically comparable spaces $\mathcal{X}$ and $\mathcal{Y}$.
@@ 3 Methodology
where we additionally enforce the approximate joint constraint. Notice the difference with Equation 4 and Equation 5: given the structure of the parametric distributions we use, it is possible to show that $\nabla_{\theta}\mathbb{E}_{x,y\sim p^{\theta}_{X,Y}}[-\log p^{\theta}_{X|Y}]\approx\nabla_{\theta}\mathbb{E}_{x,y\sim p^{\theta}_{X,Y}}[-\log p^{\phi}_{Y|Y}]$ whenever $p^{\theta}_{X,Y}=p^{\phi}_{X,Y}$. Strict adherence to the joint constraint, in principle, allows “swapping” the roles of the conditional models without affecting the optimization dynamics, leading to a cooperative method. In practice, we found trough empirical exploration that such a cooperative formulation, albeit approximate, proves to be much more stable than the original problem, and consequently decided to adopt it in our implementation, as described next.
@@ 3.1 Practical implementation
$\operatorname*{arg\,min}_{\theta}\sum_{t=1}^{T}\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I),x_{0}\sim p_{X}}\left[||\epsilon-\epsilon^{\theta}(x_{t},t)||^{2}\right].$ (7)
@@
We being by pretraining unconditional models such that $p^{\theta_{*}}_{X}(x)\approx p_{X}(x)$ and $p^{\phi_{*}}_{Y}(y)\approx p_{Y}(y)$. Then, we use $\theta_{*}$ and $\phi_{*}$ to initialize conditional models $p^{\theta}_{X|Y}$ and $p^{\phi}_{Y|X}$, which use denoising networks that accept additional conditioning signals, following Zhang et al. (2023). Next, we interpret the optimization expressed in Section 3 as a model fine-tuning objective, which is reminiscent of the work by Fan et al. (2023)
@@
$\mathbb{E}_{p^{\theta}_{X,Y}}r^{\phi}(y,x)\sum_{t=1}^{T}\nabla_{\theta}\log p^{\theta}(x_{t-1}|x_{t},y),$ (10)
@@
$\sum_{t=1}^{T}\nabla_{\theta}\mathbb{E}_{x_{t}}\left[||\epsilon^{\theta}(x_{t},y=\emptyset,t)-\epsilon^{\theta_{*}}(x_{t},t)||^{2}\right].$ (11)
@@
  Update $\theta$ using Equation 10 and Equation 11
@@ 4 Experiments
ddmec is a general method that can be applied across a variety of data domains, as it relies on an information-theoretic measure to match unpaired entities. Next, we demonstrate ddmec versatility using two realistic pairing tasks that use various data modalities, including multi-omics and image data. We compare ddmec to state-of-the-art methods for each task, and measure performance using domain-specific metrics. Details about ddmec implementation, and our experimental protocol are given in Appendix A.
@@ 4.1 Multi-omics single-cell alignment
We compare our proposed method ddmec to several baselines, both from the machine learning and the bio-informatics literature, including info-ot (Chuang et al., 2023), scot (Demetci et al., 2022), mmd-ma (Liu et al., 2019), and unioncom (Cao et al., 2020). The first two methods, scot and info-ot, propose variants of an OT formulation, the first based on the Gromov-Wasserstein distance which preserves local neighborhood geometry when moving data points, whereas the second is an information theoretic extension of OT that maximizes the mutual information between domains while minimizing geometric distances. mmd-ma is a global manifold alignment algorithm based on the maximum mean discrepancy measure, whereas unioncom performs unsupervised topological alignment for single-cell multi-omics data that emphasizes both local and global alignment. All alternative methods we consider require geometric distances or similarity measures, which is a pain point that our method ddmec lifts completely.
@@
In our experiments we use the SnaReSeq (Chen et al., 2019) dataset, which links chromatin accessibility with gene expression data on a mixture of four cells types. We use the same preprocessing procedures detailed in (Demetci et al., 2022), which deal with filtering spurious data affected by technical errors, and normalization. Data samples have 1–1 correspondence information, which constitute the groud-truth information used for our performance evaluation. We use the average “fraction of samples closer than the true match (foscttm)” metric introduced by Liu et al. (2019): given a sample in one domain, this amounts to compute the fraction of samples that are positioned more closely to it than its true match after pairing. Results report the average foscttm across all samples, where lower values indicate better performance. We also report the label transfer accuracy as done by Cao et al. (2020), which measures how well sample labels are transferred between domains based on neighborhood alignment. A $k$-nearest neighbor classifier is trained on one domain and used to predict labels in the other.
Results.
SNAREseq
FOS $\downarrow$ Acc $\uparrow$
@@ Results.
[…]
ddmec 0.147 98.6
Table 1: Single-Cell alignment experiments.
Table 1 presents the coupling results for the single-cell alignment problem, in which results marked with $\star$ are reported from (Chuang et al., 2023). We remark that ddmec outperforms alternatives in terms of the foscttm metric, and is on-par with the best method in terms of accuracy. In this experiment, ddmec is trained once, and inference is conducted five times with different seeds. Unlike other methods, ddmec is conceptually different as it generates samples rather than learning a deterministic, 1-1 mapping. To compute the different metrics, given a sample from one modality, we use ddmec to generate a coupling to the other modality, then select the nearest sample from the dataset based on Euclidean distance. Figure 1 illustrates the conditional generation of ddmec in both directions, plotted using PCA: the cell type is correctly transferred from one modality to another, as the generated samples (triangles) largely match the ground truth samples (Circles).
@@ 4.2 Unpaired image translation
As the literature on image translation is vast, here we primarily focus on the unpaired case, and compare our method to a vast range of alternatives. Generative Adversarial Network (gan) have been widely applied to this domain (Pang et al., 2021). These methods can be broadly categorized into those focusing on cycle-consistency, which enforces bidirectional mappings between image domains, such as CycleGAN (Zhu et al., 2017), DualGAN (Yi et al., 2017), SCAN (Van Gansbeke et al., 2020), and U-GAT-IT (Kim, 2019); the second category uses distance-based methods, such as DistanceGAN (Benaim & Wolf, 2017), GCGAN (Fu et al., 2019), CUT (Park et al., 2020), and LSeSim (Zheng et al., 2021). Diffusion-based models, which are related to our method, have also been explored for unpaired image translation. UNIT-DDPM (Sasaki et al., 2021) learns two conditional models along with two additional domain translation models, incorporating a GAN-like cycle-consistency loss. ILVR (Choi et al., 2021) and SDEdit (Meng et al., 2021) utilize a diffusion model in the target domain while conditioning on a source image to refine the sampling procedure for image translation. EGSDE (Zhao et al., 2022) employs an energy function pretrained on both source and target domains to guide the inference process. Similarly, SDDM (Sun et al., 2023) introduces manifold constraints, forcing distributions at adjacent time steps to be decomposable into denoising and refinement components. Compared to these methods, ddmec leverages two conditional models, one per domain, which can be initialized using pretrained unconditional diffusion models. By design, our method does not require comparable domains and does not rely on a specific image similarity measure. We report results according to two values for the guidance coefficient, a parameter influencing conditional generation.
@@ Results.
This outcome aligns with expectations, as ddmec is designed to reduce uncertainty and enforce adherence to marginal constraints. In Figure 2, qualitative results further confirm the performance of ddmec in this task. Additional results are available in Appendix B.
@@
ddmec(guidance=9) 60.70 $\pm$ 1.07 0.41 $\pm$ 0.0006
ddmec(guidance=7) 58.50 $\pm$ 0.96 0.40 $\pm$ 0.001

### v2

@@ Abstract
Machine Learning, ICML, Diffusion models, Multimodality, Unpaired data, Minimum entropy coupling
@@ 1 Introduction
Modeling multimodal data allows for a more comprehensive understanding, reflecting the inherently multi-faceted nature of the real world. Recent works in representation learning (Radford et al., 2021; Lu, 2023; Manzoor et al., 2023; Chen et al., 2023), the study of multivariate systems (McSharry et al., 2024; Liang et al., 2023; Bounoua et al., 2024b), generative modeling (Rombach et al., 2022; Tang et al., 2023b, a; Bounoua et al., 2024a; Esser et al., 2024), and multimodal conversational agents (Li et al., 2023; Liu et al., 2023; Shukor et al., 2023; Xue et al., 2024; Wu et al., 2024), are few examples to illustrate the fervent effort in the machine learning community to address and exploit multimodality. However, the intrinsic complexity of multimodal data introduces several challenges that hinder their application in machine learning research. Modality heterogeneity complicates and sometimes impedes geometric comparisons, requiring for example learning a mapping from ambient to latent spaces (Rombach et al., 2022; Tang et al., 2023b; Liu et al., 2023; Bounoua et al., 2024a) or stringent assumptions (Liang et al., 2022; Xia et al., 2023; Dong et al., 2024; Ibrahimi et al., 2024; Zhang et al., 2024b). Alignment across modalities at spatial, temporal or semantic levels is another challenge, which calls for costly pre-processing steps such as synchronization (Hanchate et al., 2024; Chen et al., 2024; Scirè, 2024; Martin-Turrero et al., 2024).
@@
In this work, we study the problem of unpaired multimodal data through the lens of coupling, a fundamental problem in probability theory, that aims at determining the optimal joint distribution of random variables given their marginal distributions, with early attempts at solving it dating back to the work by Fréchet (1951). The pairing problem belongs to a broad class of methods (Den Hollander, 2012; Lin et al., 2014; Benes & Stepán, 2012; Yu & Tan, 2018): some cast it through the lens of information-theoretic quantities, where optimality is defined in terms of Entropy minimization or Mutual Information maximization, others focus on Optimal Transport (OT) (Villani, 2009; Peyré & Cuturi, 2019), where optimality is defined as minimizing the expected value of a transport cost over the joint distribution. Our focus is the Minimum Entropy Coupling (MEC) problem, which aims at finding the joint distribution with the smallest Entropy, given the marginal distribution of some random variables. Recent applications include entropic causal inference (Kocaoglu et al., 2017; Javidian et al., 2021; Compton, 2022), communication systems (Sokota et al., 2022), steganography (de Witt et al., 2022), random number generation (Li, 2021), dimensionality reduction (Cicalese et al., 2016; Vidyasagar, 2012), lossy compression (Ebrahimi et al., 2024), and multimodal learning (Liang et al., 2024).
While the MEC problem is known to be NP-Hard (Vidyasagar, 2012; Kovačević et al., 2012), the literature contains many approximation and greedy algorithms  (Painsky et al., 2013; Kovacevic et al., 2013; Cicalese et al., 2016; Li, 2021), and theoretical studies about the approximation qualities of such approaches (Cicalese et al., 2017, 2019). Nevertheless, the vast majority of prior work on the MEC problem focus on discrete distributions: instead, we consider the continuous variant of MEC, and propose a flexible and general solution to the coupling problem for arbitrary, continuous distributions. The MEC problem for continuous random variables is much more complex than its discrete counterpart, and can be ill-defined in certain cases due to the properties of differential Entropy and the challenges inherent to continuous distributions living in an infinite dimensional space.
@@ 2 Problem Formulation
Given two random variables $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$ with marginal probability distributions $p_{X}(x)$ and $p_{Y}(y)$ respectively, we consider a parametric space $\mathcal{P}^{\theta}=\{p^{\theta}_{XY}(x,y)\}$ of joint distributions over the space $\mathcal{X}\times\mathcal{Y}$, with induced marginal distributions $p^{\theta}_{X}(x),p^{\theta}_{Y}(y)$ (where $p^{\theta}_{X}(x)\triangleq\int_{\mathcal{Y}}p^{\theta}_{XY}(x,y)\,dy$ and similarly for $p^{\theta}_{Y}(y)$). The MEC problem between the two original distributions $p_{X}(x)$ and $p_{Y}(y)$ consists in finding a joint distribution $p^{\theta}_{XY}(x,y)$ such that i) the induced marginal distributions $p^{\theta}_{X}(x),p^{\theta}_{Y}(y)$ match them either exactly or approximately and ii) the joint distribution is the one with minimal entropy (Kovacevic et al., 2013; Cicalese et al., 2017, 2019). The constraints over the search space $\mathcal{P}^{\theta}$ are referred to as marginal contraints
@@
[…]
@@ 3 Methodology
where we additionally enforce the approximate joint constraint. Notice the difference with Equation 4 and Equation 5: given the structure of the parametric distributions we use, it is possible to show that $\nabla_{\theta}\mathbb{E}_{x,y\sim p^{\theta}_{X,Y}}[-\log p^{\theta}_{X|Y}]\approx\nabla_{\theta}\mathbb{E}_{x,y\sim p^{\theta}_{X,Y}}[-\log p^{\phi}_{Y|X}]$ whenever $p^{\theta}_{X,Y}=p^{\phi}_{X,Y}$ (see Section A.1 for more details). Strict adherence to the joint constraint, in principle, allows “swapping” the roles of the conditional models without affecting the optimization dynamics, leading to a cooperative method. In practice, we found trough empirical exploration that such a cooperative formulation (See Figure 4), albeit approximate, proves to be much more stable than the original problem, and consequently decided to adopt it in our implementation, as described next.
@@ 3.1 Practical implementation
[…]
@@
We being by pretraining unconditional models such that $p^{\theta_{*}}_{X}(x)\approx p_{X}(x)$ and $p^{\phi_{*}}_{Y}(y)\approx p_{Y}(y)$. Then, we use $\theta_{*}$ and $\phi_{*}$ to initialize conditional models $p^{\theta}_{X|Y}$ and $p^{\phi}_{Y|X}$, which use denoising networks that accept additional conditioning signals, following Zhang et al. (2023). Next, we interpret the optimization expressed in Section 3 as a model fine-tuning objective, which is reminiscent of the work by Fan et al. (2023).
@@
$\mathbb{E}_{p^{\theta}_{X,Y}}r^{\phi}(y,x)\sum_{t=1}^{T}\nabla_{\theta}\log p^{\theta}(x_{t-1}|x_{t},y)$ (10)
@@
[…]
@@
 Update $\theta$ using Equations 10 and 11
@@ 4 Experiments
DDMEC  is a general method that can be applied across a variety of data domains, as it relies on an information-theoretic measure to match unpaired entities. Next, we demonstrate DDMEC  versatility using two realistic pairing tasks that use various data modalities, including multi-omics and image data. We compare DDMEC  to state-of-the-art methods for each task, and measure performance using domain-specific metrics. More details about DDMEC  implementation, and our experimental protocol are given in Appendix A.
@@ 4.1 Multi-omics single-cell alignment
We compare our proposed method DDMEC  to several baselines from both the machine learning and bioinformatics literature, including scot (Demetci et al., 2022), mmd-ma (Liu et al., 2019), unioncom (Cao et al., 2020), and scTopoGAN (Singh et al., 2023). scot proposes a variant of an OT formulation based on the Gromov-Wasserstein distance, which preserves local neighborhood geometry when transporting data points. mmd-ma is a global manifold alignment algorithm based on the maximum mean discrepancy (MMD) measure. unioncom performs unsupervised topological alignment for single-cell multi-omics data, emphasizing both local and global alignment. scTopoGAN uses topological autoencoders to obtain latent representations of each modality separately; a topology-guided Generative Adversarial Network then aligns these latent representations into a common space. We compare our method to Infoot (Chuang et al., 2023) in Section B.1. All alternative methods we consider require a choice of distance or similarity measures, which is a pain point that our method DDMEC  completely eliminates.
@@
We evaluate our method on single-cell multi-omics datasets: the peripheral blood mononuclear cells (PBMC) dataset and the bone marrow (BM) dataset. The PBMC dataset comprises healthy human peripheral blood mononuclear cells, profiled using the 10x Genomics multiome protocol, which enables simultaneous measurement of gene expression (RNA) and chromatin accessibility (ATAC) from the same cells. This dataset includes a total of 11,910 cells, encompassing 7 major immune cell types that are further subdivided into 20 finer-grained cell subclasses. The BM dataset consists of human bone marrow cells profiled using the CITE-seq protocol (Stoeckius et al., 2017), which jointly captures gene expression (RNA) and protein abundance via antibody-derived tags (ADT). A set of 10,235 cells are randomly selected from each modality based on the major cell type labels.
For both datasets, we adopt the data preprocessing and evaluation pipeline described in (Singh et al., 2023), resulting in 50-dimensional embeddings per modality. To assess the quality of the coupling, we compute cross-modal neighborhood consistency: for each cell in one modality, we identify its $k=5$ nearest neighbors in the aligned space from the other modality using Euclidean distance. We then evaluate the proportion of cases where the cell’s class or subclass label matches the majority label among its neighbors. The resulting metrics are reported as the Celltype Acc and Subcelltype Acc, respectively. In this experiment, DDMEC is trained once, and inference is conducted five times with different seeds.
Method Celltype Acc $\uparrow$ Subcelltype Acc $\uparrow$
PBMC
[…]
SCOT $\star$ 12.9 ± 1.1 2.4 ± 0.2
scTopoGAN $\star$ 61.7 ± 8.6 41.3 ± 6.5
DDMEC 66.3 ± 2.6 46.0 ± 0.5
BM
UnionCom $\star$ 51.8 ± 3.7 20.9 ± 2.6
@@
DDMEC 77.3±0.1 44.2±0.1
Table 1: Single-Cell alignment experiments
Results.
Table 1 presents the quantitative results for the single-cell alignment task. Results marked with $\star$ are reported directly from (Singh et al., 2023). We observe that DDMEC  consistently outperforms existing baselines on the PBMC dataset, achieving superior performance in aligning both coarse-grained cell types and fine-grained cell subclasses. On the BM dataset, DDMEC  obtains the best performance for subclass-level alignment and ranks second for cell-type alignment.
Notably, DDMEC  is the only method that demonstrates robust performance across both datasets, whereas alternative approaches exhibit inconsistent behavior—e.g., scot performs well on BM but fails to generalize to PBMC. In contrast to existing methods that learn deterministic one-to-one mappings between modalities, DDMEC  is fundamentally generative. It learns to sample from a coupling distribution rather than enforcing a fixed correspondence. To compute alignment metrics, we draw a sample in the target modality conditioned on a source cell, then identify the closest observed cell in the dataset using Euclidean distance. Figure 1 illustrates conditional generation with DDMEC using UMAP projections.
Figure 1: Conditional generation with DDMEC  on PBMC dataset. Top: UMAP visualizations of source and generated data. Bottom: points are colored by cell type, illustrating how well DDMEC   preserves cell type separation.
@@ 4.2 Unpaired image translation
As the literature on image translation is vast, here we primarily focus on the unpaired case, and compare our method to a vast range of alternatives. Generative Adversarial Network (gan) have been widely applied to this domain (Pang et al., 2021). These methods can be broadly categorized into those focusing on cycle-consistency, which enforces bidirectional mappings between image domains, such as CycleGAN (Zhu et al., 2017), DualGAN (Yi et al., 2017), SCAN (Van Gansbeke et al., 2020), and U-GAT-IT (Kim, 2019); the second category uses distance-based methods, such as DistanceGAN (Benaim & Wolf, 2017), GCGAN (Fu et al., 2019), CUT (Park et al., 2020), and LSeSim (Zheng et al., 2021). Diffusion-based models, which are related to our method, have also been explored for unpaired image translation. UNIT-DDPM (Sasaki et al., 2021) learns two conditional models along with two additional domain translation models, incorporating a GAN-like cycle-consistency loss. ILVR (Choi et al., 2021) and SDEdit (Meng et al., 2021) utilize a diffusion model in the target domain while conditioning on a source image to refine the sampling procedure for image translation. EGSDE (Zhao et al., 2022) employs an energy function pretrained on both source and target domains to guide the inference process. Similarly, SDDM (Sun et al., 2023) introduces manifold constraints, forcing distributions at adjacent time steps to be decomposable into denoising and refinement components. Compared to these methods, DDMEC  leverages two conditional models, one per domain, which can be initialized using pretrained unconditional diffusion models. By design, our method does not require comparable domains and does not rely on a specific image similarity measure. We report results for two values of the guidance coefficient, a parameter influencing conditional generation.
@@
In addition, due to the bidirectional architecture of DDMEC - which leverages two conditional generative models - our framework also inherently supports translation in the reverse direction. The corresponding results for the Female $\to$ Male task are presented in Appendix B.
@@ Results.
Our results on the CelebA-HQ dataset, demonstrate that DDMEC outperforms competitors on both fid and ssim even with only $50$ sampling steps. Specifically, at $50$ steps the fid improves by approximately $1$ point and the ssim by $0.02$ points, with an even greater improvement (a $3$ point fid reduction) when using $100$ sampling steps: it is well-known in the generative modeling literature that these improvements are significant. With the CelebA-HQ dataset, DDMEC benefits from a larger training set than in AFHQ animal dataset, and achieves state-of-the-art performance on image translation.
Source
Output
Source
Output
[…]
@@
[…]
50 Sampling steps
DDMEC (guidance=9) 60.70 $\pm$ 1.07 0.410 $\pm$ 0.001
DDMEC (guidance=7) 58.50 $\pm$ 0.96 0.404 $\pm$ 0.001
100 Sampling steps
DDMEC (guidance=9) 60.51 $\pm$ 1.01 0.403 $\pm$ 0.001
DDMEC (guidance=7) 57.89 $\pm$ 0.37 0.397 $\pm$ 0.001